# Advance Sharing Procedures for the Ramp Quantum Secret Sharing Schemes With the Highest Coding Rate

### v1

@@ II-D Proof of correctness
$\{(r_{1},\ldots,r_{k-L},g_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,g_{\vec{r},\vec{s}}(\alpha_{n})):\vec{r}\in\mathbf{F}_{q}^{k-L}\}.$ (6)
@@
In order to prove the theorem, it is enough to show that the sets (5) and (6) are the same. For fixed $\vec{s}\in\mathbf{F}_{q}^{L}$ and $\vec{r}\in\mathbf{F}_{q}^{k-L}$, by Proposition 2 there exists a unique polynomial $f_{\vec{c}}$ with $\vec{c}\in D(\vec{s})$. This correspondence gives a bijectition sending $\vec{r}\in\mathbf{F}_{q}^{k-L}$ to $f_{\vec{c}}$ with $\vec{c}\in D(\vec{s})$. By the definition of $g_{\vec{r},\vec{s}}$, we have
$(r_{1},\ldots,r_{k-L},g_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,g_{\vec{r},\vec{s}}(\alpha_{n}))$
@@
$=$ $(f_{\vec{c}}(\alpha_{1}),\ldots,f_{\vec{c}}(\alpha_{n})),$
@@
II-E Example of the proposed encoding
We reuse definitions from Section II-B. Let $p=2$, $q=4$, $n=3$, $k=2$, $L=1$, and $\alpha_{i}=\alpha^{i}\in\mathbf{F}_{4}$, where $\alpha$ is a primitive element of $\mathbf{F}_{4}$. Note that since $p\neq q$, this case cannot be handled in the previous research [8]. In our proposal, the dealer prepares 2-qudit entangled state $\frac{1}{2}\sum_{r\in\mathbf{F}_{4}}|r\rangle|r\rangle$, and send 1 qudit in it. Then the dealer is given a quantum secret $|\vec{s}\rangle$ for $\vec{s}=(s)\in\mathbf{F}_{4}^{1}$. Then we have
@@ II-E Example of the proposed encoding
The unitary map $U_{\mathrm{enc}}$ sends $|r\rangle|s\rangle$ to $|\frac{r-s}{\alpha_{1}}\alpha_{2}+s\rangle|\frac{r-s}{\alpha_{1}}\alpha_{3}+s\rangle$. Now it is clear that the left most share can be distributed before the secret $|s\rangle$ is given, in contrast to Section II-B.
@@ III-A Zhang and Matsumoto’s scheme
As mentioned before, Zhang and Matsumoto [14] also proposed a $(k,L,n)$ ramp quantum secret sharing scheme. Their proposal has strong security in contrast to [10], and the maximum possible number of participants was smaller, that is, $n\leq q-L$. As in Section II, $\alpha_{1}$, …, $\alpha_{n}$ are pairwise distinct elements in $\mathbf{F}_{q}$. We choose extra elements $\beta_{1}$, …, $\beta_{L}$ such that all $\alpha_{1}$, …, $\alpha_{n}$ and $\beta_{1}$, …, $\beta_{L}$ are pairwise distinct. In [14] any of $\alpha_{1}$, …, $\alpha_{n}$ and $\beta_{1}$, …, $\beta_{L}$ can be $0\in\mathbf{F}_{q}$. For $\vec{s}\in\mathbf{F}_{q}^{L}$, we define
@@ III-B Example of the original encoding
Let $p=q=7$, $n=4$, $k=3$, $L=2$, $(\alpha_{1}$, …, $\alpha_{4})=(6$, $2$, $4$, 5)5)5 ) and $(\beta_{1}$, $\beta_{2})=(1$, 3)3)3 ). This example is the same as [14, Example 2]. For $\vec{s}=(s_{1}$, $s_{2})$, $D_{\mathrm{ZM}}(\vec{s})$ contains $(c_{1}$, …, $c_{3})$ if and only if
@@
The solution of the above system of linear equations for $(c_{1}$, …, $c_{3})$ is $D_{\mathrm{ZM}}(\vec{s})=\{(5s_{1}+3s_{2}+3c_{3},3s_{1}+4s_{2}+3c_{3},c_{3}):c_{3}\in\mathbf{F}_{7}\}$, which is the same [14, Eq. (3)]. For a given quantum state $|\vec{s}\rangle=|s_{1}\rangle\otimes|s_{2}\rangle$, the quantum state of $4$ shares is (the normalizing constant $1/\sqrt{7}$ is omitted) $\sum_{c_{3}=0}^{6}\otimes_{i=1}^{4}|\alpha_{i}^{2}c_{3}+\alpha_{i}(3s_{1}+4s_{2}+3c_{3})+(5s_{1}+3s_{2}+3c_{3})\rangle=|2s_{1}+6s_{2}\rangle\otimes|4s_{1}+4s_{2}\rangle\otimes|3s_{1}+5s_{2}\rangle\otimes|6s_{1}+2s_{2}\rangle+\cdots$, which is equivalent to [14, Eq. (5)]. Since the quantum state of every qudit depends on $s_{1}$ and $s_{2}$, from the encoding procedure in [14], it seems unclear (to the authors) how one share is advance shareable.
@@ III-C Proposed advance sharing procedure for [14]
Suppose that the quantum secret is $|\vec{s}\rangle=|s_{1}\rangle\otimes\cdots\otimes|s_{L}\rangle\in\mathcal{H}_{q}^{\otimes L}$ with $\vec{s}=(s_{1}$, …, $s_{L})\in\mathbf{F}_{q}^{L}$. For $\vec{r}\in\mathbf{F}_{q}^{k-L}$ and $\vec{s}$, let $h_{\vec{r},\vec{s}}(x)$ be a univariate polynomial such that $h_{\vec{r},\vec{s}}(x)\in D_{\mathrm{ZM}}(\vec{s})$ and $h_{\vec{r},\vec{s}}(x)(\alpha_{i})=r_{i}$ for $i=1$, …, $k-L$.
@@
Let $f(x)\in\mathcal{P}_{k}$ as defined in Lemma 1. The condition $f(x)\in D_{\mathrm{ZM}}(\vec{s})$ means that $f(\beta_{i})=s_{i}$ for $i=1$, …, $L$. Together with the conditions $f(x)(\alpha_{i})=r_{i}$ for $i=1$, …, $k-L$, the required $h_{\vec{r},\vec{s}}(x)$ is a $f(x)$ such that $\mathrm{ev}^{\prime}(f)=(s_{1}$, …, $s_{L}$, $r_{1}$, …, $r_{k-L})$, where $\mathrm{ev}^{\prime}(f)=(f(\beta_{1})$, …, $f(\beta_{L})$, $f(\alpha_{1})$, …, $f(\alpha_{k-L}))$. By Lemma 1, we can see that such a polynomial $f(x)$ exists uniquely, which is $h_{\vec{r},\vec{s}}(x)$. ∎
By Proposition 5, we can define a unitary map $U_{\mathrm{ZM,enc}}$ from $\mathcal{H}_{q}^{\otimes k}$ to $\mathcal{H}_{q}^{\otimes(n-k+L)}$ sending a quantum state $|\vec{r}\rangle|\vec{s}\rangle$ to $|h_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|h_{\vec{r},\vec{s}}(\alpha_{n})\rangle$.
@@
    $\frac{1}{q^{k-L}}\sum_{\vec{r}\in\mathbf{F}_{q}^{k-L}}|\vec{r}\rangle|h_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|h_{\vec{r},\vec{s}}(\alpha_{n})\rangle\in\mathcal{H}_{q}^{\otimes n}.$ (9)
@@ III-D Proof of correctness
In this subsection, we will prove that the proposed procedure gives the same quantum state of shares as the original [14] for a given quantum secret $|\vec{s}\rangle$. In Eq. (8), the set of indices appearing in the quantum state is
@@
In order to prove the theorem, it is enough to show that the sets (10) and (11) are the same. For fixed $\vec{s}\in\mathbf{F}_{q}^{L}$ and $\vec{r}\in\mathbf{F}_{q}^{k-L}$, by Proposition 5 there exists a unique polynomial $f_{\vec{c}}(x)$ with $\vec{c}\in D_{\mathrm{ZM}}(\vec{s})$. This correspondence gives a bijectition between $\vec{r}\in\mathbf{F}_{q}^{k-L}$ to $f_{\vec{c}}(x)$ with $\vec{c}\in D(\vec{s})$. By the definition of $h_{\vec{r},\vec{s}}$, we have
@@ III-E Example of the proposed encoding
We reuse definitions from Section III-B. Let $p=q=7$, $n=4$, $k=3$, $L=2$, $(\alpha_{1}$, …, $\alpha_{4})=(6$, $2$, $4$, 5)5)5 ) and $(\beta_{1}$, $\beta_{2})=(1$, 3)3)3 ). In our proposal, the dealer prepares 2-qudit entangled state $\frac{1}{\sqrt{7}}\sum_{r\in\mathbf{F}_{7}}|r\rangle|r\rangle$, and send 1 qudit in it. Then the dealer is given a quantum secret $|\vec{s}\rangle$ for $\vec{s}=(s_{1},s_{2})\in\mathbf{F}_{7}^{2}$. Then we have
@@ IV Concluding remarks
In this paper, we showed how to make shares in Ogawa et al.’s [10] and Zhang-Matsumoto [14] ramp quantum secret sharing scheme advance shareable, and also proved that the new proposed encoding procedures retain all properties, such as coding rates, access structures, strong security from the original schemes. We also proved that the advance shareable sets are made as large as possible in the proposed encoding procedures.
Similarly to [12], there could be a general method to construct advance sharing procedure for any stabilizer-based quantum secret sharing, while retaining the correspondence between secrets and shares. Its investigation could be a future research agenda.
@@ Acknowledgment
This research is in part supported by the JSPS grant No. 23K10980.

### v2

@@ II-E Proof of correctness
$V_{2}$ $=$ $\{(r_{1},\ldots,r_{k-L},g_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,g_{\vec{r},\vec{s}}(\alpha_{n}))$ (8)
$:\vec{r}\in\mathbf{F}_{q}^{k-L}\}.$
@@
For a fixed quantum secret $|\vec{s}\rangle$ for $\vec{s}\in\mathbf{F}_{q}^{L}$, the proposed procedure gives the same quantum state of shares as the original scheme by Ogawa et al. [7].
@@
Equation (2) can be written as
$\frac{1}{\sqrt{q^{k-L}}}\sum_{(v_{1},\ldots,v_{n})\in V_{1}}|v_{1}\rangle\otimes\cdots\otimes|v_{n}\rangle,$ (9)
and Eq. (6) can be written as
$\frac{1}{\sqrt{q^{k-L}}}\sum_{(v_{1},\ldots,v_{n})\in V_{2}}|v_{1}\rangle\otimes\cdots\otimes|v_{n}\rangle.$ (10)
To prove the theorem, we have to prove that Eqs. (9) and (10) are equal. In order to prove the equality between Eqs. (9) and (10), it is enough to show that the sets (7) and (8) are the same. To show the equality between two sets (7) and (8), in Eq. (11) we will show that there is a one-to-one correspondence between elements in the two sets (7) and (8).
For fixed $\vec{s}\in\mathbf{F}_{q}^{L}$ and $\vec{r}\in\mathbf{F}_{q}^{k-L}$, by Proposition 2 there exists a unique polynomial $f_{\vec{c}}$ with $\vec{c}\in D(\vec{s})$. This correspondence gives a bijection sending $\vec{r}\in\mathbf{F}_{q}^{k-L}$ to $f_{\vec{c}}$ with $\vec{c}\in D(\vec{s})$. By the definition of $g_{\vec{r},\vec{s}}$, we have
$(r_{1},\ldots,r_{k-L},g_{\vec{r},\vec{s}}(\alpha_{k-L+1}),\ldots,g_{\vec{r},\vec{s}}(\alpha_{n}))\in V_{2}$ (11)
@@
$=$ $(f_{\vec{c}}(\alpha_{1}),\ldots,f_{\vec{c}}(\alpha_{n}))\in V_{1},$
@@
II-F Example of the proposed encoding
We reuse definitions from Section II-C. Let $q=4$, $n=3$, $k=2$, $L=1$, and $\alpha_{i}=\alpha^{i}\in\mathbf{F}_{4}$, where $\alpha$ is a primitive element of $\mathbf{F}_{4}$. Note that since $q$ is not a prime number, this case cannot be handled in the previous research [14]. In our proposal, the dealer prepares 2-qudit entangled state $\frac{1}{2}\sum_{r\in\mathbf{F}_{4}}|r\rangle|r\rangle$, and send 1 qudit in it. Then the dealer is given a quantum secret $|\vec{s}\rangle$ for $\vec{s}=(s)\in\mathbf{F}_{4}^{1}$. Then we have
@@ II-F Example of the proposed encoding
The unitary map $U_{\mathrm{enc}}$ sends $|r\rangle|s\rangle$ to $|\frac{r-s}{\alpha_{1}}\alpha_{2}+s\rangle|\frac{r-s}{\alpha_{1}}\alpha_{3}+s\rangle$. Now it is clear that the left most share can be distributed before the secret $|s\rangle$ is given, in contrast to Section II-C.
@@ III-A Zhang and Matsumoto’s scheme
As mentioned before, Zhang and Matsumoto [10] also proposed a $(k,L,n)$ ramp quantum secret sharing scheme. Their proposal has strong security in contrast to Ogawa et al. [7], and the maximum possible number of participants was smaller, that is, $n\leq q-L$. As in Section II, $\alpha_{1}$, …, $\alpha_{n}$ are pairwise distinct elements in $\mathbf{F}_{q}$. We choose extra elements $\beta_{1}$, …, $\beta_{L}$ such that all $\alpha_{1}$, …, $\alpha_{n}$ and $\beta_{1}$, …, $\beta_{L}$ are pairwise distinct. In [10] any of $\alpha_{1}$, …, $\alpha_{n}$ and $\beta_{1}$, …, $\beta_{L}$ can be $0\in\mathbf{F}_{q}$. For $\vec{s}\in\mathbf{F}_{q}^{L}$, we define
@@ III-B Example of the original encoding
Let $q=7$, $n=4$, $k=3$, $L=2$, $(\alpha_{1}$, …, $\alpha_{4})=(6$, $2$, $4$, 5)5)5 ) and $(\beta_{1}$, $\beta_{2})=(1$, 3)3)3 ). This example is the same as [10, Example 2]. For $\vec{s}=(s_{1}$, $s_{2})$, $D_{\mathrm{ZM}}(\vec{s})$ contains $(c_{1}$, …, $c_{3})$ if and only if
@@
The solution of the above system of linear equations for $(c_{1}$, …, $c_{3})$ is $D_{\mathrm{ZM}}(\vec{s})=\{(5s_{1}+3s_{2}+3c_{3},3s_{1}+4s_{2}+3c_{3},c_{3}):c_{3}\in\mathbf{F}_{7}\}$, which is the same [10, Eq. (3)]. For a given quantum state $|\vec{s}\rangle=|s_{1}\rangle\otimes|s_{2}\rangle$, the quantum state of $4$ shares is (the normalizing constant $1/\sqrt{7}$ is omitted) $\sum_{c_{3}=0}^{6}\otimes_{i=1}^{4}|\alpha_{i}^{2}c_{3}+\alpha_{i}(3s_{1}+4s_{2}+3c_{3})+(5s_{1}+3s_{2}+3c_{3})\rangle=|2s_{1}+6s_{2}\rangle\otimes|4s_{1}+4s_{2}\rangle\otimes|3s_{1}+5s_{2}\rangle\otimes|6s_{1}+2s_{2}\rangle+\cdots$, which is equivalent to [10, Eq. (5)]. Since the quantum state of every qudit seems to depend on $s_{1}$ and $s_{2}$, from the encoding procedure in [10], it seems unclear how one share is advance shareable.
@@ III-C Proposed advance sharing procedure for [10]
Suppose that the quantum secret is $|\vec{s}\rangle=|s_{1}\rangle\otimes\cdots\otimes|s_{L}\rangle\in\mathcal{H}_{q}^{\otimes L}$ with $\vec{s}=(s_{1}$, …, $s_{L})\in\mathbf{F}_{q}^{L}$. For $\vec{r}\in\mathbf{F}_{q}^{k-L}$ and $\vec{s}$, let $h_{\vec{r},\vec{s}}(x)$ be a univariate polynomial such that $h_{\vec{r},\vec{s}}(x)\in\{f_{\vec{c}}(x)$ : $\vec{c}\in D_{\mathrm{ZM}}(\vec{s})\}$ and $h_{\vec{r},\vec{s}}(x)(\alpha_{i})=r_{i}$ for $i=1$, …, $k-L$.
@@
Let $\mathcal{P}_{k}$ as defined in Lemma 1, and consider a polynomial $f(x)\in\mathcal{P}_{k}$. The condition $f(x)\in\{f_{\vec{c}}(x)$ : $\vec{c}\in D_{\mathrm{ZM}}(\vec{s})\}$ means that $f(\beta_{i})=s_{i}$ for $i=1$, …, $L$. Together with the conditions $f(\alpha_{i})=r_{i}$ for $i=1$, …, $k-L$, the required $h_{\vec{r},\vec{s}}(x)$ is a polynomial $f(x)\in\mathcal{P}_{k}$ such that $\mathrm{ev}^{\prime}(f)=(s_{1}$, …, $s_{L}$, $r_{1}$, …, $r_{k-L})$, where $\mathrm{ev}^{\prime}(f)=(f(\beta_{1})$, …, $f(\beta_{L})$, $f(\alpha_{1})$, …, $f(\alpha_{k-L}))$. By Lemma 1, we can see that such a polynomial $f(x)$ exists uniquely, which is $h_{\vec{r},\vec{s}}(x)$. ∎
Remark 7
Similarly to Remark 3, we will give a more explict description of $h_{\vec{r},\vec{s}}(x)$, which is determined by the equality condition $(h_{\vec{r},\vec{s}}(\beta_{1})$, …, $h_{\vec{r},\vec{s}}(\beta_{L})$, $h_{\vec{r},\vec{s}}(\alpha_{1})$, …, $h_{\vec{r},\vec{s}}(\alpha_{k-L}))=(s_{1}$, …, $s_{L}$, $r_{1}$, …, $r_{k-L})$. The determination of $h_{\vec{r},\vec{s}}(x)$ is known as the Lagrange interpolation [15, Chapter 13], which can be computed as follows. In order to use the Lagrange interpolation formula, define $\beta_{L+j}=\alpha_{j}$ for $j=1$, …, $k-L$, and
$\ell_{i}(x)=\prod_{1\leq j\leq k,j\neq i}\frac{x-\beta_{j}}{\beta_{i}-\beta_{j}}.$
Now $h_{\vec{r},\vec{s}}(x)$ can be obtained as
$h_{\vec{r},\vec{s}}(x)=\sum_{i=1}^{L}s_{i}\ell_{i}(x)+\sum_{i=1}^{k-L}r_{i}\ell_{i+L}(x).$
Let us describe a new encoding procedure enabling advance sharing. By Proposition 6, we can define a unitary map $U_{\mathrm{ZM,enc}}$ from $\mathcal{H}_{q}^{\otimes k}$ to $\mathcal{H}_{q}^{\otimes(n-k+L)}$ sending a quantum state $|\vec{r}\rangle|\vec{s}\rangle$ to $|h_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|h_{\vec{r},\vec{s}}(\alpha_{n})\rangle$. Note that $k=n-k+L$ by [7, Lemma 2].
@@
    $\frac{1}{\sqrt{q^{k-L}}}\sum_{\vec{r}\in\mathbf{F}_{q}^{k-L}}|\vec{r}\rangle\otimes|h_{\vec{r},\vec{s}}(\alpha_{k-L+1})\rangle\otimes\cdots\otimes|h_{\vec{r},\vec{s}}(\alpha_{n})\rangle.$ (14)
@@ III-D Proof of correctness
By the same reason explained at the beginning of Section II-E, in this subsection, we will prove that the proposed procedure gives the same quantum state of shares as the original [10] for a given quantum secret $|\vec{s}\rangle$. In Eq. (13), the set of indices appearing in the quantum state is
@@
In order to prove the theorem, it is enough to show that the sets (15) and (16) are the same. For fixed $\vec{s}\in\mathbf{F}_{q}^{L}$ and $\vec{r}\in\mathbf{F}_{q}^{k-L}$, by Proposition 6 there exists a unique polynomial $f_{\vec{c}}(x)$ with $\vec{c}\in D_{\mathrm{ZM}}(\vec{s})$. This correspondence gives a bijection between $\vec{r}\in\mathbf{F}_{q}^{k-L}$ to $f_{\vec{c}}(x)$ with $\vec{c}\in D(\vec{s})$. By the definition of $h_{\vec{r},\vec{s}}$, we have
@@ III-E Example of the proposed encoding
We reuse definitions from Section III-B. Let $q=7$, $n=4$, $k=3$, $L=2$, $(\alpha_{1}$, …, $\alpha_{4})=(6$, $2$, $4$, 5)5)5 ) and $(\beta_{1}$, $\beta_{2})=(1$, 3)3)3 ). In our proposal, the dealer prepares 2-qudit entangled state $\frac{1}{\sqrt{7}}\sum_{r\in\mathbf{F}_{7}}|r\rangle|r\rangle$, and send 1 qudit in it. Then the dealer is given a quantum secret $|\vec{s}\rangle$ for $\vec{s}=(s_{1},s_{2})\in\mathbf{F}_{7}^{2}$. Then we have
@@ IV Concluding remarks
In this paper, we proposed new encoding procedures for Ogawa et al.’s [7] and Zhang-Matsumoto [10] ramp quantum secret sharing schemes, and allowed the dealer to distribute shares in those schemes before secrets are given. This enhances applicability of the secret sharing schemes by Ogawa et al. [7] and by Zhang and Matsumoto [10] to wider scenarios in which some participants are unavailable when the dealer obtains a secret. We also proved that the new proposed encoding procedures retain desirable properties, such as coding rates, access structures, strong security from the original schemes. We also proved that the advance shareable sets are made as large as possible in the proposed encoding procedures.
Similarly to [13], there could exist a general method to construct advance sharing procedure for any stabilizer-based quantum secret sharing, while retaining the correspondence between secrets and shares. Its investigation could be a future research agenda.
As mentioned in Section II-A, it seems also important to study advance sharing with other scenarios of quantum secret sharing, for example, [18, 19, 20]. It is also another research agenda.
@@ Acknowledgment
The author would like to deeply thank reviewers’ helpful comments that improved this paper.